1) A flashlight has 6 batteries, 2 of which are defective.

a) If 2 are selected at random without replacement, find the probability that both are defective.

b) If 2 are selected at random with replacement, find the probability that both are defective.

2) How many different ways can an instructor select 4 textbooks from a possible 17?
Answer #1

Solution:

( 1 )

total number batteries = 6

number of defective batteries = 2

probability of selecting a defective battery = 2 / 6 = 0.3333

( a )

If 2 are selected at random without replacement, the probability that both are defective = ( 2/6 )*(1/5)

= 0.0666

( b )

If 2 are selected at random with replacement,the probability that both are defective.= ( 2/6 )* (2/6)

= 0.1111

( 2 )

total number of books = 17

number of textbooks to be selected = 4

number of different ways of selecting 4 books out of 17 books = 17C4 = 2380
Answer #2
1/3*1/5=1/15